### v2

@@ 1 Introduction
To address these challenges, explainability in AI typically operates at three levels: pre-model, focusing on data-level insights (e.g., feature engineering to highlight key biomarkers); in-model, leveraging inherently interpretable models (e.g., decision trees or linear models); and post-model, generating explanations after predictions (e.g., heatmaps or saliency maps). While in-model approaches often struggle to capture complex patterns, post-model XAI tools, such as Grad-CAM [5], have gained significant attraction due to their applicability to deep learning models. However, such heatmap-based methods face three key challenges in medical imaging, particularly in Alzheimer’s disease: (i) the absence of ground truth for validating visual explanations against actual dementia-related regions; (ii) insufficient quantitative metrics for evaluating alignment with known pathology; and (iii) the neglect of pre-model strategies. Techniques like voxel-based morphometry (VBM), which identify structural brain changes through statistical methods, are often overlooked in training pipelines, leading to potentially suboptimal performance. Notably, there is a dearth of effective pre-model approaches for highlighting meaningful neuroanatomical changes before model training, which could enhance interpretability.
This paper introduces Jacobian Maps (JMs) as a pre-model approach to VBM for enhancing explainability in AD detection. By computing Jacobian determinants across an image, JMs generate a matrix that accurately captures spatial and directional changes in medical scans, serving as a subject-specific ground truth. By pinpointing voxel-level cerebral changes relative to a healthy brain, JMs highlight morphometric variations, improving both interpretability and trust in AD detection. Our main contributions are:
@@
  • Improved AD detection without model modification: As a pre-model preprocessing step, JM enhances detection accuracy across all AD stages without altering the AI model. Our approach achieves 95.2% (CN), 96.3% (MCI), 90.2% (MLD), and 90.2% (SEV), outperforming traditional methods (88.3%, 90.5%, 83.4%, and 83.4%, respectively).
@@ 3.2 Brain Image Processing
Data Collection. We utilize 1557 MRI images from the OASIS-3 dataset [21], which comprises 1377 participants: 755 cognitively normal adults and 622 with varying cognitive decline, aged 42–95. To avoid overlap between training and test sets, images from the same subject are confined to one dataset partition, ensuring the evaluation of unseen data for accurate generalization. Clinical Dementia Rating (CDR) scores (0 to 3) guide the diagnosis, where 0 indicates no dementia, and 3 indicates severe dementia. To address the issue of limited data for severely demented cases, we merge moderate and severe stages into one category, resulting in four groups: normal (CN), mild cognitive impairment (MCI), mild dementia (MLD), and moderate-to-severe (SEV). This classification provides more granularity than binary methods (AD and non-AD).
@@ 4.1 Model and Training
Figure 3: Performance comparison per each AD class between conventional registration (without JM) versus registration with JM (gray-scale vs red-scale bars) using three metrics.
@@ 4.2 Model Performance Evaluation
Figure 3 compares the model’s performance using metrics such as accuracy, precision, and recall across registered images and JMs. The cross-validation, applied across AD classes (Normal, MCI, Mild, and Moderate), shows that JMs consistently outperform registered images in all metrics. These results underscore the effectiveness of JMs in capturing more informative patterns for AD classification, leading to improved model performance and diagnostic accuracy.
Figures 4 further illustrate the learning curves for models trained on registered images versus JMs. Validation accuracies are noticeably more consistent with JMs, indicating greater stability and reliability throughout training. Additionally, both training and validation performances improve with JMs, confirming the benefits of their enhanced feature representation.
@@ 4.3 Interoperability Evaluation
The rankings in Table 1 highlight which brain regions contribute most significantly to distinguishing between brain states (Normal, MCI, Mild, and Moderate). Several key observations can be made:
@@
    • This suggests that patterns in the frontal and temporal areas are crucial for differentiating the normal brain state.
@@
    • This shift implies that changes in the Temporal Lobe play a larger role in the early stages of cognitive decline.
@@
    • This progression underscores the increasing prominence of the Temporal Lobe during this stage of Alzheimer’s disease.
@@
    • This broader distribution of importance suggests more widespread neurodegeneration, involving frontal and parietal regions as the disease advances.
These findings highlight the dynamic changes in brain region importance as Alzheimer’s disease progresses from Normal to MCI, Mild, and Moderate states. While some regions, such as the Frontal-Temporal and Temporal Lobe, remain critical throughout, others, such as the Frontal and Parietal Lobes, gain prominence in later stages.
@@ 4.4 Extension to Multi-modal Setting
We employ the same model architecture outlined in Section 4.1, with an adaptation of two additional convolutional layers to optimize computational efficiency in handling the increased dimensionality from multimodal fusion. This adjustment ensures that the model effectively captures the richer feature representations offered by both MRI and CT data.
@@
Table 2: Multimodal: ablation study for Registration vs. JM
@@
    CN - Cognitively Normal, MCI - Mild Cognitive Impairment, SEV - Severe Dementia, REG - Registered without JM.

### v3

@@ 1 Introduction
To address these challenges, explainability in AI typically operates at three levels: pre-model, focusing on data-level insights (e.g., feature engineering to highlight key biomarkers); in-model, leveraging inherently interpretable models (e.g., decision trees or linear models); and post-model, generating explanations after predictions (e.g., heatmaps or saliency maps). While in-model approaches often struggle to capture complex patterns, post-model XAI tools, such as Grad-CAM [5], have gained significant attraction due to their applicability to deep learning models. However, such heatmap-based methods face key challenges in medical imaging, particularly in Alzheimer’s disease: (i) the absence of ground truth for validating visual explanations against actual dementia-related regions; (ii) insufficient quantitative metrics for evaluating alignment with known pathology. Techniques like voxel-based morphometry (VBM), which identify structural brain changes through statistical methods, are often overlooked in training pipelines, leading to potentially suboptimal performance. Notably, there is a dearth of effective pre-model approaches for highlighting meaningful neuroanatomical changes before model training, which could enhance interpretability.
This paper introduces Jacobian Maps (JMs), derived from the VBM preprocessing pipeline but repurposed in this paper as a pre-model anatomical reference, to enhance explainability in AD detection. By computing Jacobian determinants across an image, JMs generate a matrix that accurately captures spatial and directional changes in medical scans, serving as a subject-specific ground truth. By pinpointing voxel-level cerebral changes relative to a healthy brain, JMs highlight morphometric variations, improving both interpretability and trust in AD detection. Our main contributions are:
@@
  • Improved AD detection without model modification: As a pre-model preprocessing step, JM enhances detection accuracy across all AD stages without altering the AI model. Our approach achieves 95.2% (CN), 96.3% (MCI), 90.2% (MLD), and 90.2% (MOD), outperforming traditional methods (88.3%, 90.5%, 83.4%, and 83.4%, respectively).
@@ 3.2 Brain Image Processing
Data Collection. We utilize 1557 MRI images from the OASIS-3 dataset [21], which comprises 1377 participants: 755 cognitively normal adults and 622 with varying cognitive decline, aged 42–95. To avoid overlap between training and test sets, images from the same subject are confined to one dataset partition, ensuring the evaluation of unseen data for accurate generalization. Clinical Dementia Rating (CDR) scores (0 to 3) guide the diagnosis, where 0 indicates no dementia, and 3 indicates severe dementia. To address the issue of limited data for severely demented cases, we merge moderate and severe stages into one category, resulting in four groups: normal (CN), mild cognitive impairment (MCI), mild dementia (MLD), and moderate-to-severe (MOD). This classification provides more granularity than binary methods (AD and non-AD).
@@ 4.1 Model and Training
Figure 3: Performance comparison per each AD class between conventional registration (without JM; gray-scale bars) versus registration with JM (red-scale bars).
@@ 4.2 Model Performance Evaluation
Fig. 3 compares the model’s performance using metrics such as accuracy, precision, and recall across registered images and JMs. The cross-validation, applied across AD classes (Normal, MCI, Mild, and Moderate), shows that JMs consistently outperform registered images in all metrics. These results underscore the effectiveness of JMs in capturing informative local volumetric changes (expansion/compression) in brain anatomy, which enhances model’s discriminative power and leads to improved model performance and diagnostic accuracy.
Fig. 4 further illustrate the learning curves for models trained on registered images versus JMs. Models trained on JM exhibit smoother learning curves and stabler validation accuracy compared to those trained on conventionally registered images. The top row (no JM) shows greater fluctuations, while the bottom row (with JM) demonstrates more consistent convergence in addition to improved training and validation performance, highlighting the benefits of JM’s enhanced feature representation.
@@ 4.3 Interoperability Evaluation
The rankings in Table 1 highlight which brain regions contribute most significantly to distinguishing between brain states (Normal, MCI, Mild-Moderate, and Severe). Several key observations can be made:
@@
    • The Frontal-temporal region is involved in memory, decision-making and language, and is known to degenerate early in AD. This explains its highest rank across all stages.
@@
    • The increasing prominence of the Temporal lobe during the MCI and MLD stages is because this lobe includes hippocampus and entorhinal cortex, which are among the first regions to show atrophy, with memory loss being a key symptom.
@@
    • The Frontal and Parietal Lobes are responsible for planning, judgment, spatial orientation, and attention. Our finding shows that these areas are less impacted in early stages yet become more prominently affected in later stages as neurodegeneration spreads.
These findings highlight the dynamic changes in brain region importance as Alzheimer’s disease progresses from Normal to MCI, Mild, and Moderate states. Importantly, they are consistent with clinical evidence, which underscores the significance and validity of our XAI method.
@@ 4.4 Extension to Multi-modal Setting
We employ the same model architecture outlined in Section 4.1, with an adaptation of two additional convolutional layers to optimize computational efficiency in handling the increased dimensionality from multimodal fusion. This adjustment allows the model to more effectively capture the richer feature representations offered by both MRI and CT data.
@@
Table 2: Multimodal: Ablation for REG (without JM) vs. JM
@@
    CN - Cognitively Normal, MCI - Mild Cognitive Impairment, MOD - Moderate-to-severe Dementia, REG - Registered without JM.